OpenStudy (anonymous):

Use Lagrange multipliers to find minimum and maximum values of the function subject to the constraints given: f(x,y,z)=(x^2)*y+z subject to x^2+y^2=1, z=y

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

I can't solve the equations after differentiating the components. dL/dx=2xy+2xr=0 dL/dy=x^2+2yr=0 dL/dz=x^2+1+2zr=0 so from the first equation: 2xy+2xr=0 x(y+r)=0 ; therefore x=0, y=-r second equation: x^2+2yr=0 0+2yr=0 2(-r)r=0 ; therefore r=0 third equation: x^2+1+2zr=0 0+1+2z(0)=0 1=0 <--- I guess I'm stuck here

OpenStudy (zarkon):

you have 2 constraints...are you using two Lagrange multipliers?

OpenStudy (anonymous):

One constraint function: z=y or g(y,z)=y-z and one function that needs to be optimized: f(x,y,z)=(x^2)*y+z subject to x^2+y^2=1